# Equations involving the modular $j$-function and its derivatives

### v1

@@ 6. Zero Estimates
On the other hand, by Proposition 6.4, on a Zariski open dense set of fundamental domains, the denominator has order at most $(-(s+t+\ell),0)$ at the cusp, while the numerator has order at most $(-\deg(p),0)$, thus whenever $h_{\ell^{\prime}}\neq 0$ we also have
@@ 7.1. $\mathbf{j}$-Homogeneous equations
The $\mathbf{j}$-degree of $F$ is the degree of $F(X,Y_{0},T^{2}Y_{1},T^{4}Y_{2})$ in $T$. We say that $F$ is $\mathbf{j}$-homogeneous if $F(X,Y_{0},T^{2}Y_{1},T^{4}Y_{2})$ is homogeneous in the variable $T$.
@@
First, we observe that
@@
where $h(X,W)=X^{4}+2\frac{j^{\prime}(W)}{j^{\prime\prime}(W)}X^{3}$. If we can find $\tau\in\operatorname{\mathbb{H}}$ such that $j^{\prime\prime}(\tau)=0\neq j^{\prime}(\tau)$, we are done by Proposition 5.4, so suppose by contradiction that this does not happen. In this case, $\nicefrac{{j^{\prime}(z)}}{{j^{\prime\prime}(z)}}$ is bounded on the standard fundamental domain $\mathbb{F}$: by construction, it cannot have poles in $\operatorname{\mathbb{H}}$, and by looking at the $q$-expansions, it is also bounded for $\operatorname{Im}(z)\to+\infty$. This also implies that $h(\tau+\nicefrac{{d}}{{c}},\tau)\neq 0$ for any $\nicefrac{{d}}{{c}}\in\operatorname{\mathbb{Q}}$, $\tau\in\operatorname{\mathbb{H}}$ except possibly when $j^{\prime}(\tau)=0$.
Second, under the above assumptions, we shall verify that $h(\tau+r,\tau)\neq 0$ for all $r\in\operatorname{\mathbb{R}}$, $\tau\in\operatorname{\mathbb{H}}$ (7.3.1), and in turn deduce that
@@
for all $z\in\mathbb{F}$, for some $\varepsilon>0$ (7.3.2).
@@
where $M$ is a bound for $\left|\frac{j^{\prime}(z)}{j^{\prime\prime}(z)}\right|$ on $\mathbb{F}$, and so $\frac{j^{\prime}(z)}{j^{\prime\prime}(z)}\to 0$ as $\operatorname{Im}(z)\to 0$. By the Schwarz Reflection Principle, $\frac{j^{\prime}(z)}{j^{\prime\prime}(z)}$ extends to a holomorphic function on $\operatorname{\mathbb{C}}$ that vanishes for all $z\in\operatorname{\mathbb{R}}$, and is thus constantly 0, a contradiction.
For a general $\mathbf{j}$-homogeneous $F$, we just need to find an appropriate generalisation of Equation 7.1 and fill the details in the above sketch.
@@
Let $F\in\operatorname{\mathbb{C}}[\mathbf{Y}]$ be $\mathbf{j}$-homogeneous. Then there are polynomials $p_{k}\in\operatorname{\mathbb{C}}[\mathbf{Y}]$ such that for all $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ with $c\neq 0$ we have
@@
Let $F$ be as in the hypothesis and fix the polynomials $p_{k}$, $h$ as in the conclusion of Lemma 7.2.
If some $\frac{p_{k}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ has a pole in $\operatorname{\mathbb{H}}$ or exponential growth in some fundamental domain, we are done by Proposition 5.4. Therefore, we shall assume that this is not the case. In particular, we assume that $h(z+\nicefrac{{d}}{{c}},\mathbf{j}(z))$ has no pole in $\operatorname{\mathbb{H}}$ for any $\nicefrac{{d}}{{c}}$.
With this additional assumption, if $F(\mathbf{j}(\tau))=0$, then $F(\mathbf{j}(\gamma\tau))=0$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. Since $F$ is not divisible by $Y_{0}-j(\tau)$, Proposition 6.2 implies that $j^{\prime}(\tau)=0$. In particular, for any $\nicefrac{{d}}{{c}}$, $h(\tau+\nicefrac{{d}}{{c}},\mathbf{j}(\tau))=0$ implies that $j^{\prime}(\tau)=0$.
@@
For every $u\in\operatorname{\mathbb{R}}$, the function $h(z+u,\mathbf{j}(z))$ has no zero in $\operatorname{\mathbb{H}}$.
@@
Since $h$ is monic in $Z$, the polynomial $h(\tau+Z,\mathbf{j}(\tau))$ is not identically zero, thus for $r$ sufficiently close to $u$ we have $h(\tau+r,\mathbf{j}(\tau))\neq 0$, and so $\tau_{r}\neq\tau$. Therefore, the $\tau_{r}$’s can be chosen to accumulate at $\tau$ for $r\to u$. However, our assumptions imply that $j^{\prime}(\tau_{r})=0$, so the $\tau_{r}$’s lie in a discrete subset of $\operatorname{\mathbb{H}}$ (namely, the orbits of $\rho$ and $i$), a contradiction. ∎
@@
Our current assumptions imply that each $\frac{p_{k}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ has no pole in $\overline{\mathbb{F}}$ nor exponential growth in $\mathbb{F}$, and so is bounded in $\overline{\mathbb{F}}$, say by $M>0$. In particular, since $h$ is monic in $Z$, we have $|h(z+u,\mathbf{j}(z))|>\frac{1}{2}|z+u|^{N}$ as soon as $|z+u|>2N\!M$. On the other hand, $|z+u|\leq 2N\!M$ defines a compact subset $K$ of $\overline{\mathbb{F}}\times\operatorname{\mathbb{R}}$, and so the function
@@
attains some minimum $\varepsilon^{\prime}>0$ on $K$, since it does not vanish by 7.3.1. The conclusion follows on taking $\varepsilon=\min\{\varepsilon^{\prime},\frac{1}{2}\}$. ∎
@@
Therefore, $\frac{p_{k_{0}}(\mathbf{j}(z))}{F(\mathbf{j}(z))}\to 0$ as $\operatorname{Im}(z)\to 0$, thus by Schwarz’s reflection principle, $\frac{p_{k_{0}}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ has a holomorphic extension to $\operatorname{\mathbb{C}}$ that is constantly zero on $\operatorname{\mathbb{R}}$, thus constantly zero on $\operatorname{\mathbb{C}}$, hence $p_{k_{0}}=0$, a contradiction. ∎
@@
Let $S=\{z\in\operatorname{\mathbb{H}}:F(\mathbf{j}(z))=0\}$ for some $F$ as in Theorem 7.3. As observed in the proof, every $\tau\in S$ that is not in the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbit of $\rho$ or $i$ must also be a pole of some coefficient $\frac{p_{k}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ appearing in Lemma 7.2. Combining this information with Proposition 4.1, it follows that $j(S)$ is infinite and every point of $j(S)\setminus\{0,1728\}$ is an accumulation point of $j(S)$.
@@
Given different polynomials $\alpha_{1},\ldots,\alpha_{\ell}\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$, we use $\alpha_{1}^{\mathbb{N}}\cdots\alpha_{\ell}^{\mathbb{N}}$ to denote the set of all products between elements of different $\alpha_{i}^{\mathbb{N}}$.
@@
Immediate. ∎
@@
Immediate after the change of variables $C=\frac{Y_{2}}{2Y_{1}}$. ∎
@@
We recall that $F^{\gamma}=F^{\Gamma}(a,b,c,d,X,\mathbf{Y})$ for any $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. If $F\in\operatorname{\mathbb{C}}[Y_{0}]$, the conclusion follows by Corollary 6.3, so we may assume that this is not the case, and in particular that $n+N>0$.
We observe immediately that $h_{n+N}=C^{n}p_{N}(\nicefrac{{A}}{{C}},\mathbf{Y})$. Moreover,
$h_{0}=D^{n}\sum_{k=0}^{N}p_{k}(\nicefrac{{B}}{{D}},C,\mathbf{Y})D^{k}=\sum_{k=\operatorname{ord_{\mathbf{j}}}(F)}^{N}p_{k}(\nicefrac{{B}}{{D}},C,\mathbf{Y})D^{k}.$
@@
has a Zariski dense set of solutions (by Corollary 5.5 in the first case, and Proposition 5.4 in the second one), hence so does $F(z,\mathbf{j}(z))=0$ (by Proposition 5.1), as desired.
@@
Suppose that $p_{N}$ is not in $Y_{0}^{\mathbb{N}}(Y_{0}-1728)^{\mathbb{N}}Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[Z]$. Recall that $p_{N}$ is $\mathbf{j}$-homogeneous by Proposition 7.5. Since $h_{n+N}=C^{n}p_{N}(\nicefrac{{A}}{{C}},\mathbf{Y})$, if $p_{N}$ depends on $Y_{2}$, so does $h_{n+N}(\nicefrac{{a}}{{c}},\mathbf{Y})$ for all but finitely many values of $\nicefrac{{a}}{{c}}$.
If $p_{N}$ does not depend on $Y_{2}$, then $p_{N}=Y_{1}^{N}h(\nicefrac{{A}}{{C}},Y_{0})$ for some polynomial $h\in\operatorname{\mathbb{C}}[Z,Y_{0}]$, and by assumption $h(\nicefrac{{A}}{{C}},Y_{0})$ has at least one root distinct from 00 and $1728$, seen as a polynomial in $Y_{0}$ (in an algebraic closure of $\operatorname{\mathbb{C}}(\nicefrac{{A}}{{C}})$), in which case so does $h(\nicefrac{{a}}{{c}},Y_{0})$ except for finitely many values of $\nicefrac{{a}}{{c}}$.
In either case, since the map $\gamma\mapsto\nicefrac{{a}}{{c}}$ from $\operatorname{SL}_{2}(\operatorname{\mathbb{C}})$ to $\operatorname{\mathbb{C}}$ is dominant, we get that for all $\gamma$ except on some proper Zariski closed subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$, the polynomial $h_{n+N}(a,b,c,d,\mathbf{Y})=c^{n}p_{N}(\nicefrac{{a}}{{c}},\mathbf{Y})$ has an irreducible factor $r\in\operatorname{\mathbb{C}}[\mathbf{Y}]$ such that $r(\mathbf{j}(z))=0$ has a Zariski dense set of solutions (by respectively Theorem 7.3, after noticing that the factors of a $\mathbf{j}$-homogeneous polynomial are $\mathbf{j}$-homogeneous, and Corollary 6.3).
Suppose that $F$ is in $\operatorname{\mathbb{C}}[X,Y_{0}]$. By irreducibility of $F$, we have that $F=F(X,Y_{0})$ is not divisible by $Y_{0}$ nor $Y_{0}-1728$. In particular, $p_{N}=\Gamma(F)=F(Z,Y_{0})$ is also not divisible by $Y_{0}$ nor $Y_{0}-1728$, which puts us back in the previous case.
Suppose that $p_{N}$ is in $Y_{0}^{\mathbb{N}}(Y_{0}-1728)^{\mathbb{N}}Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[Z]$ and $F$ is not in $\operatorname{\mathbb{C}}[X,Y_{0}]$. Write $p_{N}=r(Z)Y_{0}^{s}(Y_{0}-1728)^{t}Y_{1}^{\ell}$. Since $F$ is irreducible, $F$ is also not in $Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[X,Y_{0}]$, thus by Corollary 7.8,
@@
in which case the function $\frac{p_{\operatorname{ord_{\mathbf{j}}}(F)}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{p_{N}(\nicefrac{{a}}{{c}},\mathbf{j}(z))}$ has a pole at some $\tau\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$ or exponential growth in some fundamental domain by Corollary 6.8 (recall that by Proposition 7.7 we know that $p_{\operatorname{ord_{\mathbf{j}}}(F)}$ does not depend on $Y_{2}$). If we fix some $\nicefrac{{b}}{{d}}$, $c$ as above, and a corresponding pole $\tau$ or a fundamental domain $\eta\mathbb{F}$ with exponential growth, then the function
$\frac{h_{0}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{h_{n+N}(\nicefrac{{a}}{{c}},\mathbf{j}(z))}=\frac{1}{c^{n}r(\nicefrac{{a}}{{c}})}\sum_{k=\operatorname{ord_{\mathbf{j}}}(F)}^{N}\frac{p_{k}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{j(z)^{s}(j(z)-1728)^{t}j^{\prime}(z)^{\ell}}d^{k},$
has no pole at $\tau$ nor exponential growth in $\eta\mathbb{F}$ only when $d$ satisfies a non-trivial polynomial equation over $\nicefrac{{b}}{{d}}$, $c$. Since $\gamma\mapsto(\nicefrac{{b}}{{d}},c,d)$ is also a dominant map, for $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ the above function has a pole at $\tau$ or exponential growth in $\eta\mathbb{F}$, as claimed. ∎

### v3

@@ 6. Zero Estimates
On the other hand, by Proposition˜6.4, on a Zariski open dense set of fundamental domains, the denominator has order at most $(-(s+t+\ell),0)$ at the cusp, while the numerator has order at most $(-\deg(p),0)$, thus whenever $p_{\ell^{\prime}}\neq 0$ we also have
@@ 7.1. $\mathbf{j}$-Homogeneous equations
The $\mathbf{j}$-degree of $F\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ is the degree of $F(X,Y_{0},T^{2}Y_{1},T^{4}Y_{2})$ in $T$, which we denote $\deg_{\mathbf{j}}(F)$. We say that $F$ is $\mathbf{j}$-homogeneous if $F(X,Y_{0},T^{2}Y_{1},T^{4}Y_{2})$ is homogeneous in the variable $T$.
@@
First, we observe that for $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{SL}_{2}(\operatorname{\mathbb{Z}})$,
@@
[…]
Second, under the above assumptions, we shall verify that $h(\tau+r,\tau)\neq 0$ for all $r\in\operatorname{\mathbb{R}}$, $\tau\in\operatorname{\mathbb{H}}$ (˜7.3.1), and in turn deduce that
@@
for all $z\in\mathbb{F}$, for some $\varepsilon>0$ (˜7.3.2).
@@
[…]
For a general $\mathbf{j}$-homogeneous $F$, we just need to find an appropriate generalisation of Equation˜7.1 and fill the details in the above sketch.
@@
Let $F\in\operatorname{\mathbb{C}}[\mathbf{Y}]$ be $\mathbf{j}$-homogeneous. Then there are polynomials $p_{k}\in\operatorname{\mathbb{C}}[\mathbf{Y}]$ and $h\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$ such that for all $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ with $c\neq 0$ we have
@@
Let $F$ be as in the hypothesis and fix the polynomials $p_{k}$, $h$ as in the conclusion of Lemma˜7.2.
[…]
With this additional assumption, if $F(\mathbf{j}(\tau))=0$, then $F(\mathbf{j}(\gamma\tau))=0$ for all $\gamma\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. Since $F$ is not divisible by $Y_{0}-j(\tau)$, Proposition˜6.2 implies that $j^{\prime}(\tau)=0$. In particular, for any $\nicefrac{{d}}{{c}}$, $h(\tau+\nicefrac{{d}}{{c}},\mathbf{j}(\tau))=0$ implies that $j^{\prime}(\tau)=0$.
@@
[…]
@@
[…]
@@
Our current assumptions imply that each $\frac{p_{k}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ has neither a pole in $\overline{\mathbb{F}}$ nor exponential growth in $\mathbb{F}$. Since these functions are in $\mathcal{P}$ rather than $\mathcal{P}(w)$, their order at $i\infty$ is of the form $(e,0)$, thus at least $(0,0)$, and so they are bounded in $\overline{\mathbb{F}}$, say by $M>0$. In particular, since $h$ is monic in $Z$, we have $|h(z+u,\mathbf{j}(z))|>\frac{1}{2}|z+u|^{N}$ as soon as |z+u|>2NM|z+u|>2N\!M. On the other hand, $|z+u|\leq 2N\!M$ defines a compact subset $K$ of $\overline{\mathbb{F}}\times\operatorname{\mathbb{R}}$, and so the function
@@
attains some minimum $\varepsilon^{\prime}>0$ on $K$, since it does not vanish by ˜7.3.1. The conclusion follows on taking $\varepsilon=\min\{\varepsilon^{\prime},\frac{1}{2}\}$. ∎
@@
[…]
@@
Let $S=\{z\in\operatorname{\mathbb{H}}:F(\mathbf{j}(z))=0\}$ for some $F$ as in Theorem˜7.3. As observed in the proof, every $\tau\in S$ that is not in the $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$-orbit of $\rho$ or $i$ must also be a pole of some coefficient $\frac{p_{k}(\mathbf{j}(z))}{F(\mathbf{j}(z))}$ appearing in Lemma˜7.2. Combining this information with Proposition˜4.1, it follows that $j(S)$ is infinite and every point of $j(S)\setminus\{0,1728\}$ is an accumulation point of $j(S)$.
@@
Given different polynomials $\alpha_{1},\ldots,\alpha_{\ell}\in\operatorname{\mathbb{C}}[X,\mathbf{Y}]$, we use $\alpha_{1}^{\mathbb{N}}\cdots\alpha_{\ell}^{\mathbb{N}}$ to denote the set of all products between elements of different $\alpha_{t}^{\mathbb{N}}$.
@@
Let $X^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}}$ denote a monomial, so $\alpha,\beta_{0},\beta_{1},\beta_{2}\in\mathbb{N}$. Observe that
$\Gamma(X^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}})=Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}W^{2\beta_{1}+3\beta_{2}}(WY_{2}+2CY_{1})^{\beta_{2}}.$
Hence
$\deg_{W}(\Gamma(X^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}}))=2\beta_{1}+4\beta_{2}$
and the term accompanying $W^{2\beta_{1}+4\beta_{2}}$ is $Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}}$.
Now write $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\beta_{2})$ and
$F(X,\mathbf{Y})=\sum_{(\alpha,\boldsymbol{\beta})\in\mathbb{N}^{4}}c_{\alpha,\boldsymbol{\beta}}X^{\alpha}\mathbf{Y}^{\boldsymbol{\beta}},$
where $\mathbf{Y}^{\boldsymbol{\beta}}=Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}}$ and $c_{\alpha,\boldsymbol{\beta}}\in\operatorname{\mathbb{C}}$. Then, since $\Gamma$ is a homomorphism by (O1),
$\Gamma(F)=\sum_{(\alpha,\boldsymbol{\beta})\in\mathbb{N}^{4}}c_{\alpha,\boldsymbol{\beta}}\Gamma\left(X^{\alpha}\mathbf{Y}^{\boldsymbol{\beta}}\right).$
Since $N=\deg_{W}(\Gamma(F))$, then
$p_{N}=\sum_{(\alpha,\boldsymbol{\beta})\in\mathbb{N}^{4}:2\beta_{1}+4\beta_{2}=N}c_{\alpha,\boldsymbol{\beta}}Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}}.$
From this we see that $p_{N}$ does not depend on $C$ and that $p_{N}$ is $\mathbf{j}$-homogeneous. This expression also gives us that $p_{N}$ is the coefficient of $W^{N}$ in $F(Z,Y_{0},W^{2}Y_{1},W^{4}Y_{2})$, and so $N=\deg_{W}(F(Z,Y_{0},W^{2}Y_{1},W^{4}Y_{2}))=\deg_{\mathbf{j}}(F)$. ∎
@@
We proceed as in the proof of Proposition˜7.5. From
$\Gamma(X^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}Y_{2}^{\beta_{2}})=Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}}W^{2\beta_{1}+3\beta_{2}}(WY_{2}+2CY_{1})^{\beta_{2}}$
we see that the smallest power of $W$ appearing in this expression is $2\beta_{1}+3\beta_{2}$, and it is accompanied by
$2^{\beta_{2}}C^{\beta_{2}}Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}+\beta_{2}},$
which does not depend on $Y_{2}$. Hence
$p_{k_{0}}=\sum_{(\alpha,\boldsymbol{\beta})\in\mathbb{N}^{4}:2\beta_{1}+3\beta_{2}=k_{0}}c_{\alpha,\boldsymbol{\beta}}2^{\beta_{2}}C^{\beta_{2}}Z^{\alpha}Y_{0}^{\beta_{0}}Y_{1}^{\beta_{1}+\beta_{2}},$
which also shows that $p_{k_{0}}$ is the coefficient of $W^{k_{0}}$ in $F(Z,Y_{0},W^{2}Y_{1},2CW^{3}Y_{1})$. Using the change of variables $C=\frac{Y_{2}}{2Y_{1}}$, we conclude that $k_{0}=\deg_{\mathbf{j}}(F)$, and since $k_{0}=2\beta_{1}+3\beta_{2}\geq 2(\beta_{1}+\beta_{2})$ we have $k_{0}\geq 2\deg_{Y_{1}}(p_{k_{0}})$, concluding the proof. ∎
@@
We recall that $F^{\gamma}=F^{\Gamma}(a,b,c,d,X,\mathbf{Y})$ for any $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$. If $F\in\operatorname{\mathbb{C}}[Y_{0}]$, the conclusion follows by Corollary˜6.3, so we may assume that this is not the case, and in particular that $n+N>0$.
We observe immediately that $h_{n+N}=C^{n}p_{N}(\nicefrac{{A}}{{C}},\mathbf{Y})$ using Proposition˜7.5. Moreover, by Proposition˜7.7,
$h_{0}=D^{n}\sum_{k=0}^{N}p_{k}(\nicefrac{{B}}{{D}},C,\mathbf{Y})D^{k}=D^{n}\sum_{k=\operatorname{ord_{\mathbf{j}}}(F)}^{N}p_{k}(\nicefrac{{B}}{{D}},C,\mathbf{Y})D^{k}.$
@@
has a Zariski dense set of solutions (by Corollary˜5.5 in the first case, and Proposition˜5.4 in the second one), hence so does $F(z,\mathbf{j}(z))=0$ (by Proposition 5.1), as desired.
@@
Suppose that $p_{N}$ is not in $Y_{0}^{\mathbb{N}}(Y_{0}-1728)^{\mathbb{N}}Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[Z]$. Recall that $p_{N}$ is $\mathbf{j}$-homogeneous by Proposition˜7.5. Since $h_{n+N}=C^{n}p_{N}(\nicefrac{{A}}{{C}},\mathbf{Y})$, if $p_{N}$ depends on $Y_{2}$, so does $h_{n+N}(\nicefrac{{a}}{{c}},\mathbf{Y})$ for all but finitely many values of $\nicefrac{{a}}{{c}}$.
If $p_{N}$ does not depend on $Y_{2}$, then $p_{N}=Y_{1}^{N}h(\nicefrac{{A}}{{C}},Y_{0})$ for some polynomial $h\in\operatorname{\mathbb{C}}[Z,Y_{0}]$, and by assumption $h(\nicefrac{{A}}{{C}},Y_{0})$ has at least one root distinct from 0 and $1728$, seen as a polynomial in $Y_{0}$ (in an algebraic closure of $\operatorname{\mathbb{C}}(\nicefrac{{A}}{{C}})$), in which case so does $h(\nicefrac{{a}}{{c}},Y_{0})$ except for finitely many values of $\nicefrac{{a}}{{c}}$.
[…]
Suppose that $F$ is in $\operatorname{\mathbb{C}}[X,Y_{0}]$. By irreducibility of $F$, we have that $F=F(X,Y_{0})$ is neither divisible by $Y_{0}$ nor $Y_{0}-1728$. In particular, $p_{N}=\Gamma(F)=F(Z,Y_{0})$ is also neither divisible by $Y_{0}$ nor $Y_{0}-1728$, which puts us back in the previous case.
Suppose that $p_{N}$ is in $Y_{0}^{\mathbb{N}}(Y_{0}-1728)^{\mathbb{N}}Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[Z]$ and $F$ is not in $\operatorname{\mathbb{C}}[X,Y_{0}]$. Write $p_{N}=r(Z)Y_{0}^{s}(Y_{0}-1728)^{t}Y_{1}^{\ell}$. Since $F$ is irreducible, $F$ is also not in $Y_{1}^{\mathbb{N}}\operatorname{\mathbb{C}}[X,Y_{0}]$, thus by Corollary˜7.8,
@@
in which case the function $\frac{p_{\operatorname{ord_{\mathbf{j}}}(F)}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{p_{N}(\nicefrac{{a}}{{c}},\mathbf{j}(z))}$ has a pole at some $\tau\in\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})\rho\cup\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})i$ or exponential growth in some fundamental domain by Corollary˜6.8 (recall that by Proposition 7.7 we know that $p_{\operatorname{ord_{\mathbf{j}}}(F)}$ does not depend on $Y_{2}$). If we fix some $\nicefrac{{b}}{{d}}$, $c$ as above, and a corresponding pole $\tau$ or a fundamental domain $\eta\mathbb{F}$ with exponential growth, then the function
$\frac{h_{0}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{h_{n+N}(\nicefrac{{a}}{{c}},\mathbf{j}(z))}=\frac{d^{n}}{c^{n}r(\nicefrac{{a}}{{c}})}\sum_{k=\operatorname{ord_{\mathbf{j}}}(F)}^{N}\frac{p_{k}(\nicefrac{{b}}{{d}},c,\mathbf{j}(z))}{j(z)^{s}(j(z)-1728)^{t}j^{\prime}(z)^{\ell}}d^{k},$
has neither a pole at $\tau$ nor exponential growth in $\eta\mathbb{F}$ only when $d$ satisfies a non-trivial polynomial equation over $\nicefrac{{b}}{{d}}$, $c$. Since $\gamma\mapsto(\nicefrac{{b}}{{d}},c,d)$ is also a dominant map, for $\gamma$ in a Zariski open dense subset of $\operatorname{SL}_{2}(\operatorname{\mathbb{Z}})$ the above function has a pole at $\tau$ or exponential growth in $\eta\mathbb{F}$, as claimed. ∎